# UAV-Assisted Resilience in 6G and Beyond Network Energy Saving: A Multi-Agent DRL Approach

### v1

@@ I Introduction
The rapid proliferation of mobile devices and and the shift toward 6G and beyond networks have driven an unprecedented surge in energy consumption. This escalating energy demand not only increases operational expenditure for mobile network operators but also amplifies the environmental impact through higher carbon emissions. Consequently, improving the energy efficiency (EE) and sustainability of wireless infrastructures has become a central design goal for next-generation networks. Beyond energy concerns, modern communication systems must remain resilient to failures caused by natural disasters, backhaul congestion, or equipment malfunctions, necessitating rapid and reliable service recovery under dynamic conditions [2].
@@ II System Model
We considers a downlink communication network in which UAVs provide service to ground users when nearby cells are temporarily turned off to save energy or due to power outages or failures. The goal is to simultaneously optimize the UAV trajectory and transmit power to maximize user coverage while minimizing the UAV’s energy consumption.
As shown in Fig. 1, each GBS consists of three sectors or cells. Each cell serves the associated mobile users within its coverage area. When the cells encounter problems such as hardware failures, natural disasters, or energy-saving problems, but the traffic demand from mobile users in cells suddenly increases unexpectedly, turning on the cells when they are in turn-off/deep sleep modes will take many minutes. Even when turned on, these cells are still overloaded and cannot serve all users in time. Therefore, we use UAVs acting as ABSs to serve users in a timely manner. We design the scheduling so that users are allocated to the UAV with the best SINR. This ensures that each user is connected to the most favorable serving UAV based on the instantaneous channel quality and the UAV’s location.
@@ II-1 Cells
There are $K$ fixed cells, each is a hexagonal cell. The cell operational state is given by $s_{k}(t)\in\{0,1\}$, where 1 denotes ON and 0 denotes OFF. The cell ON/OFF schedule is predetermined during data synthesizing process and provided as input. The set of inactive cells requiring UAV assistance at time $t$ is: $A(t)=\{k|s_{k}(t)=0\}$.
@@ II-A Communication Model
Given the location $(x_{j},\ y_{j},\ 0)$ of each user $j\in M$ at ground level and the current UAV location $(x_{i},\ y_{i},\ H)$ in time slot n, the distance $d_{ij}$ between the UAV and the user is:
$d_{ij}(t)=\sqrt{H^{2}+(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}}.$ (4)
The channel between the user $j$ and the UAV $i$ follows a Rician fading channel model with a factor K, where the channel coefficient $h_{ij}$ can be written as:
$h_{ij}(t)=\sqrt{w_{ij}(t)}.\tilde{h}_{ij}(t),$ (5)
where $w_{ij}(t)$ represents for the large-scale fading effects and $h_{ij}(t)$ accounts for Rician small-scale fading coefficient. Specifically, $w_{ij}(t)$ can be modeled as:
$w_{ij}(t)=w_{0}.d_{ij}^{-\alpha}(t),$ (6)
where $w_{0}$ is the average channel power gain at the reference distance $d=1m$, and $\alpha\geq 2$ is the path loss exponent for the Rician fading channel. The small scale fading $\tilde{h}_{ij}$ with an expected value $\mathop{\mathbb{E}}[|\tilde{h}_{ij}^{2}|]=1$, is given by:
$\tilde{h}_{ij}(t)=\sqrt{\frac{G}{1+G}}\bar{h}_{ij}(t)\ +\sqrt{\frac{1}{1+G}}\hat{h}_{ij}(t),$ (7)
where $\bar{h}_{ij}(t)$ and $\hat{h}_{ij}(t)\sim\mathcal{CN}(0,\,1)$ denote the deterministic LoS and non-line-of-sight (NLoS) component (Rayleigh fading) during time slot t, respectively; G is the Rician factor. The SINR of each user $j$ is:
$\gamma_{ij}(t)=\frac{P_{i}(t).|h_{ij}(t)|^{2}}{\sigma^{2}+\sum_{k\neq i}P_{k}(t).|h_{kj}(t)|^{2}}\ ,$ (8)
where $P_{i}(t)$ is the power of the UAV $i$ at the timeslot t, $\sigma^{2}$ is the thermal noise power which is linearly proportional to the allocated bandwidth, interference includes other UAVs and active GBSs. Thus, the data rate achievable for user $j$ when served by UAV $i$ at time $t$ is:
@@
where $B$ is the frequency range allocated to this communication link.
Each user is associated with the UAV that provides the highest SINR:
$a_{ij}(t)=\begin{cases}1,&\text{if }i=\displaystyle\arg\max_{k}\gamma_{kj}(t),\\ 0,&\text{otherwise.}\end{cases}$
Each user can be served by at most one UAV:
$\sum_{i}a_{ij}(t)\leq 1,\quad\forall j.$ (10)
To ensure satisfactory communication quality, each user must achieve a minimum data rate requirement
$R_{ij}(t)\geq R_{\mathrm{req},j},\quad\forall j$ (11)
where $R_{\mathrm{req},j}$ denotes the required data rate threshold of user $j$.
The total network throughput at time $t$ is expressed as
$R_{\mathrm{total}}(t)=\sum_{i}\sum_{j}a_{ij}(t)R_{ij}(t).$ (12)
The objective of the UAV communication system is typically to maximize $R_{\mathrm{total}}(t)$ under the above user assignment and QoS constraints.
@@ II-D Problem Formulation
The optimization objective of maximizing the users coverage while minimizing the total UAVs’ energy consumption by optimizing the UAVs’ trajectory, UAVs’ transmit power, and user-UAV association. Our problem is formulated as:
$\mathcal{P}_{1}:$ (17)
$\max_{\{q_{i}(t),P_{i}(t),a_{ij}(t)\}}\sum_{t=1}^{T}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{ij}(t)R_{ij}(t)-\lambda\sum_{t=1}^{T}\sum_{i=1}^{N}E_{i}(t)$ (18)
s.t. $R_{ij}(t)\geq R_{\mathrm{req},j},\quad\forall i,j,t,$ (18a)
$\sum_{i}a_{ij}(t)\leq 1,\quad\forall j,t,$ (18b)
$a_{ij}(t)\in\{0,1\},\quad\forall i,j,t,$ (18c)
@@
$0\leq P_{i}(t)\leq P_{max},\quad\forall i,t,$ (18e)
$\sum_{i}P_{i}(t)\leq P_{max},\quad\forall t,$ (18f)
where $\lambda$ is a trade-off coefficient balancing coverage maximization and energy minimization. Constraint (18a) ensures that each served user meets its minimum QoS requirement, while (18b) guarantees that each user is associated with at most one UAV. Constraints (18d)-(18f) enforce UAV mobility and power limitations.
@@ III POMDP Formulation for UAV-Assisted Resilience in 6G NES
III-A POMDP Formulation
We model the control problem as a partially observable Markov decision process (POMDP) [6]. Each UAV $i$ observes a local state $s_{i}(t)$ while a global state $s_{global}(t)$ is available for centralized training.
III-A1 Local State of UAV $i$
$s_{i}(t)$ $=[q_{i}(t),P_{i}(t),q_{\text{other}_{i}}(t),P_{\text{other}_{i}}(t),$ (19)
${(x_{j},y_{j},R_{req,j})}_{j\in U_{i}(t)},{s_{k}(t)}_{k=1}^{K}],$
where $q_{i}(t)$ is position of UAV $i$ at time $t$; $P_{i}(t)$ is transmit power of UAV $i$ at timestep $t$; $q_{\text{other}_{i}}(t),P_{\text{other}_{i}}(t)$ are positions and powers of other UAVs; ${(x_{j},y_{j},R_{req,j})}_{j\in U_{i}(t)}$ is information of 6 nearest users within UAV $i$ coverage diameter; ${s_{k}(t)}_{k=1}^{K}$ is states of all cells.
@@ III-A2 Global State
$s_{\text{global}}(t)=[\{q_{i}(t),P_{i}(t)\}_{i=1}^{N},\{{(x_{j},y_{j},R_{req,j})}\}_{j=1}^{M},s_{k}(t)_{k=1}^{K}].$ (20)
III-A3 Action of UAV $i$
$a_{i}(t)=[\Delta x_{i}(t),\Delta y_{i}(t),P_{i}(t)].$ (21)
III-B Reward Function
The reward emphasizes coverage maximization and energy minimization, with coverage rate is measured based on successfully served UAV-needed users:
$R_{i}(t)=\omega_{1}\cdot\mathcal{R}^{C}_{i}(t)+\omega_{2}\cdot(1-\mathcal{P}_{i}^{E}(t)),$ (22)
where $\mathcal{R}^{C}_{i}(t)$ and $\mathcal{P}_{i}^{E}(t)$ are calculated, respectively, as
$\mathcal{R}^{C}_{i}(t)=\frac{|\{j:R_{ij}(t)\geq R_{\text{req},j},\,s_{\text{cell},j}(t)=0\}|}{|\{j:s_{\text{cell},j}(t)=0\}|},$ (23)
$\mathcal{P}_{i}^{E}(t)=\frac{E_{i}(t)}{E_{max}},$ (24)
with $E_{max}=P_{max}\Delta t+\alpha_{1}v_{max}^{2}+\alpha_{2}v_{max}$.
III-C Objective
Each UAV learns a policy $\pi_{i}(a_{i}|s_{i})$ that maximizes its expected cumulative discounted reward:
$\max_{\pi_{1},\pi_{2}}E\left[\sum_{t=0}^{T}\gamma^{t}R_{i}(t)\right].$ (25)
@@ IV-A CTDE-MADDPG Framework
We adopt a Centralized Training with Decentralized Execution (CTDE) variant of MADDPG for cooperative UAVs control with better memory and training efficiency as shown 2. Each UAV is modeled as an agent with its own actor–critic pair. During training, each agent’s critic has access to the global state $s_{\text{global}}(t)$ and actions of all agents, enabling coordinated learning. However, during execution, each UAV only uses its local observations $s_{i}(t)$ through the actor network, making the system scalable and robust to partial observability.
@@ IV-C2 Exploration Strategy
Besides dynamically adjusting exploration through PER, we also add temporally correlated Ornstein-Uhlenbeck (OU) Noise to actions during training to encourage stable exploration: $N_{t}=N_{t-1}+\theta(\mu-N{t-1})\Delta t+\sigma W_{t}$ with noise parameter $\sigma$ is decayed by 0.9999 per step with a minium of 0.01 [7]. After substantial number of episodes, exploration noise is disabled to exploit the learned policy. This could prevents erratic movements that could violate physical constraints, and produces smooth, correlated actions suitable for UAV dynamics.
@@ IV-C3 Interference Management
The system accounts for inter-UAVs interference when computing SINR values. Each UAV’s transmission contributes to the interference perceived by users connected to other UAVs. This coupling is reflected in the environment dynamics and achievable rate calculation. The mechanism encourages UAV agents to learn cooperative spatial and power control strategies that balance coverage maximization and energy minimmization.
@@ V-A Reward-based Analysis
Fig. 3 illustrates the convergence of the proposed MADDPG framework. The x-axis represents the Training Episodes (up to 25,000), and the y-axis is the normalized average episodic reward (per step). The average reward rises sharply during the initial exploration phase and stabilizes around 0.032. It exhibits a clear stepped improvement around the 5,000-episode mark, which coincides with the learning-rate decay. This jump is followed by a brief period of high-variance fine-tuning. The policy then settles into its final converged state of approximately 0.039, and the stable plateau from roughly 8,000 episodes onward confirms robust convergence.
@@
As shown in Fig. 4, the learned MADDPG policy maintains the highest average reward across all timesteps, outperforming both the k-nearest neighbors (KNN)-Fixed and Random baselines. The Random policy quickly collapses toward zero reward, while the KNN-Fixed method remains myopic and less adaptive. In contrast, the MADDPG agents, through cooperative learning, appears to find a more refined balance in spatial positioning and power control. This ability to make better sequential decisions yields a consistent, albeit small, performance gain throughout the episode.
Figure 4: omparison of instantaneous reward per step, averaged over 15 testing episodes.
@@ V-B Coverage Performance Evaluation
Figure 5: Comparison of Total Coverage Ratio per Episode

### v2

@@ I Introduction
The rapid proliferation of mobile devices and the shift toward 6G and beyond networks have driven an unprecedented surge in energy consumption. This escalating energy demand not only increases operational expenditure for mobile network operators but also amplifies the environmental impact through higher carbon emissions. Consequently, improving the energy efficiency (EE) and sustainability of wireless infrastructures has become a central design goal for next-generation networks. Beyond energy concerns, modern communication systems must remain resilient to failures caused by natural disasters, backhaul congestion, or equipment malfunctions, necessitating rapid and reliable service recovery under dynamic conditions [2].
@@ II System Model
We consider a downlink communication network in which UAVs provide service to ground users when nearby cells are temporarily turned off to save energy or due to power outages or failures. The goal is to simultaneously optimize the UAV trajectory and transmit power to maximize user coverage while minimizing the UAV’s energy consumption.
As shown in Fig. 1, each GBS consists of three sectors or cells. Each cell serves the associated mobile users within its coverage area. When the cells encounter problems such as hardware failures, natural disasters, or energy-saving problems, but the traffic demand from mobile users in cells suddenly increases unexpectedly, turning on the cells when they are in turn-off/deep sleep modes will take many minutes. Even when turned on, these cells are still overloaded and cannot serve all users in time. Therefore, we use UAVs acting as ABSs to serve users in a timely manner. We design the scheduling so that users are allocated to the UAV with the best signal-to-interference-plus-noise ratio (SINR). This ensures that each user is connected to the most favorable serving UAV based on the instantaneous channel quality and the UAV’s location.
@@ II-1 Cells
There are $K$ fixed cells, each is a hexagonal cell. The operational state of each cell $k$, denoted by $s_{k}(t)\in\{0,1\}$, is determined by a traffic-aware energy saving strategy. Specifically, to minimize the substantial static power consumption of GBSs, a cell is switched to sleep mode ($s_{k}(t)=0$) if the aggregated traffic demand of its associated users falls below a predefined energy-efficiency threshold $\gamma_{th}$. Conversely, the cell remains active ($s_{k}(t)=1$) during peak loads. This pre-optimized sleeping schedule serves as the baseline environment input for the subsequent UAV resilience optimization.
@@ II-A Communication Model
Given the location $(x_{j},\ y_{j},\ 0)$ of each user $j\in\mathcal{M}$ at ground level and the current UAV location $(x_{i},\ y_{i},\ H)$ at timestep $t$, the 3D Euclidean distance $d_{ij}(t)$ between UAV $i$ and user $j$ is:
$d_{ij}(t)=\sqrt{H^{2}+(x_{i}(t)-x_{j})^{2}+(y_{i}(t)-y_{j})^{2}}.$ (4)
The channel coefficient between user $j$ and UAV $i$ follows a Rician fading model:
$h_{ij}(t)=\sqrt{w_{ij}(t)}\tilde{h}_{ij}(t),$ (5)
where $w_{ij}(t)$ represents large-scale path loss:
$w_{ij}(t)=w_{0}d_{ij}^{-\alpha}(t),$ (6)
with $w_{0}$ being the reference path loss at d=1d=1m and $\alpha\geq 2$ the path loss exponent. The small-scale Rician fading coefficient $\tilde{h}_{ij}(t)$ is given by:
$\tilde{h}_{ij}(t)=\sqrt{\frac{K}{1+K}}\bar{h}_{ij}(t)+\sqrt{\frac{1}{1+K}}\hat{h}_{ij}(t),$ (7)
where $\bar{h}_{ij}(t)$ and $\hat{h}_{ij}(t)\sim\mathcal{CN}(0,\,1)$ denote the deterministic line-of-sight (LoS) and non-line-of-sight (NLoS) Rayleigh fading components, respectively, and $K$ is the Rician K-factor.
SINR experienced by user $j$ when served by UAV $i$ is:
$\gamma_{ij}(t)=\frac{P_{i}(t)|h_{ij}(t)|^{2}}{\sigma^{2}+\sum_{k\neq i}^{N}P_{k}(t)|h_{kj}(t)|^{2}+\sum_{\ell=1}^{K}s_{\ell}(t)P_{\text{BS}}|h_{\ell j}(t)|^{2}},$ (8)
where $P_{i}(t)$ is the transmit power of UAV $i$, $\sigma^{2}$ is the thermal noise power, the first summation accounts for inter-UAV interference from other active UAVs, and the second summation accounts for interference from active ground base stations (GBSs), with $s_{\ell}(t)\in\{0,1\}$ indicating whether cell $\ell$ is active and $P_{\text{BS}}$ being the GBS transmit power.
The achievable data rate for user $j$ served by UAV $i$ is:
@@
where $B$ is the allocated bandwidth.
User-UAV Association: Each user is associated with the UAV providing the highest SINR among all active UAVs:
$a_{ij}(t)=\begin{cases}1,&\text{if }i=\displaystyle\arg\max_{k\in\mathcal{N}}\gamma_{kj}(t)\text{ and }s_{\text{cell},j}(t)=0,\\ 0,&\text{otherwise,}\end{cases}$ (10)
where $s_{\text{cell},j}(t)=0$ indicates that user $j$’s home cell is in sleep mode (requiring UAV service). Each user can be served by at most one UAV:
$\sum_{i=1}^{N}a_{ij}(t)\leq 1,\quad\forall j\in\mathcal{M}.$ (11)
QoS Constraint: To ensure satisfactory service quality, each served user must achieve a minimum data rate:
$R_{ij}(t)\geq R_{\mathrm{req},j}(t),\quad\forall j:a_{ij}(t)=1,$ (12)
where $R_{\mathrm{req},j}(t)$ is the time-varying rate requirement of user $j$, synthesized based on real-world 5G traffic patterns from [15].
@@ II-D Problem Formulation
The optimization objective is to maximize user coverage (percentage of served users meeting QoS) while minimizing total UAV energy consumption over a finite horizon $T$. Formally:
$\mathcal{P}_{1}:$ $\max_{\{\mathbf{v}_{i}(t),P_{i}(t)\}_{i,t}}\quad\sum_{t=1}^{T}\omega_{1}\cdot C(t)-\omega_{2}\cdot\frac{E_{\text{UAV}}(t)}{E_{\max}}$ (17)
s.t. $R_{ij}(t)\geq R_{\mathrm{req},j},\quad\forall i,t,j,$ (17a)
$\sum_{i}a_{ij}(t)\leq 1,\quad\forall t,j,$ (18)
$a_{ij}(t)\in\{0,1\},\quad\forall i,t,j$ (19)
@@
$0\leq P_{i}(t)\leq P_{max},\quad\forall i,t,$ (19b)
$\sum_{i}P_{i}(t)\leq P_{max},\quad\forall t,$ (19c)
where $\lambda$ is a trade-off coefficient balancing coverage maximization and energy minimization. Constraint (17a) ensures that each served user meets its minimum QoS requirement, while (18) guarantees that each user is associated with at most one UAV. Constraints (19a)-(19c) enforce UAV mobility and power limitations.
@@ III POMDP Formulation for UAV-Assisted Resilience in 6G NES
We model the multi-UAV control problem as a partially observable Markov decision process (POMDP) [6] with CTDE. During training, a global state $s_{\text{global}}(t)$ is available to all agents’ critics. During execution, each UAV $i$ operates based solely on its local observation $o_{i}(t)$.
III-A State and Observation Spaces
III-A1 Local Observation of UAV $i$
Each UAV $i$ constructs a local observation vector comprising:
$o_{i}(t)=\Big[$ $\underbrace{\mathbf{q}_{i}(t),P_{i}(t)}_{\text{own state (3D)}},\underbrace{\{\mathbf{q}_{k}(t),P_{k}(t)\}_{k\neq i}}_{\text{partner UAV states}},$
$\underbrace{\{x_{j},y_{j},R_{\text{req},j}(t)\}_{j\in\mathcal{U}_{i}(t)}}_{\text{5 nearest users}},\underbrace{\{s_{k}(t)\}_{k=1}^{K}}_{\text{cell states}}\Big],$ (20)
where $\mathcal{U}_{i}(t)\subset\mathcal{M}$ denotes the set of the 5 nearest users to UAV $i$ (within a radius threshold), limited to reduce observation dimensionality.
@@ III-A2 Global State
For centralized training, the global state aggregates information about all users:
$s_{\text{global}}(t)=\Big[\{x_{j},y_{j},R_{\text{req},j}(t)\}_{j=1}^{M},\{s_{k}(t)\}_{k=1}^{K}\Big]$ (21)
III-B Action Space
The continuous action of UAV $i$ at time $t$ is:
$\mathbf{a}_{i}(t)=[v_{x,i}(t),v_{y,i}(t),P_{i}(t)]\in\mathbb{R}^{3},$ (22)
where:
  • $v_{x,i}(t),v_{y,i}(t)\in[-v_{\max},v_{\max}]$: velocity components (m/s),
  • $P_{i}(t)\in[0,P_{\max}]$: transmit power (Watts).
III-C Reward Function
The per-timestep reward for UAV $i$ balances coverage contribution and energy penalty:
$r_{i}(t)=\omega_{1}\cdot\mathcal{C}_{i}(t)-\omega_{2}\cdot\frac{E_{i}(t)}{E_{\max}},$ (23)
where $\mathcal{C}_{i}(t)$, $E_{i}(t)$ and $E_{max}$ are calculated, respectively, as:
  • $\mathcal{C}_{i}(t)=\frac{|\{j:a_{ij}(t)=1,R_{ij}(t)\geq R_{\text{req},j}(t)\}|}{|\{j:s_{\text{cell},j}(t)=0\}|}$,
  • $E_{i}(t)=P_{i}(t)\Delta t+\alpha_{1}\|\mathbf{v}_{i}(t)\|^{2}+\alpha_{2}\|\mathbf{v}_{i}(t)\|$,
  • $E_{\max}=P_{max}\Delta t+\alpha_{1}\|\mathbf{v}_{max}(t)\|^{2}+\alpha_{2}\|\mathbf{v}_{max}(t)\|$
III-D Learning Objective
Each UAV learns a stochastic policy $\pi_{i}(\mathbf{a}_{i}|o_{i})$ to maximize expected cumulative discounted return:
$J(\pi_{1},\ldots,\pi_{N})=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{T}\gamma^{t}r_{i}(t)\right],$ (24)
where $\gamma$ is the discount factor and $\tau$ denotes a trajectory.
@@ IV-A CTDE-MADDPG Framework
We adopt a CTDE variant of MADDPG for cooperative UAVs control with better memory and training efficiency as shown 2. Each UAV is modeled as an agent with its own actor–critic pair. During training, each agent’s critic has access to the global state $s_{\text{global}}(t)$ and actions of all agents, enabling coordinated learning. However, during execution, each UAV only uses its local observations $s_{i}(t)$ through the actor network, making the system scalable and robust to partial observability.
@@ IV-C2 Exploration Strategy
Besides dynamically adjusting exploration through PER, we also add temporally correlated Ornstein-Uhlenbeck (OU) Noise to actions during training to encourage stable exploration: $N_{t}=N_{t-1}+\theta(\mu-N{t-1})\Delta t+\sigma W_{t}$ with noise parameter $\sigma$ is decayed by 0.9999 per step with a minimum of 0.01 [7]. After substantial number of episodes, exploration noise is disabled to exploit the learned policy. This prevents erratic movements that could violate physical constraints, and produces smooth, correlated actions suitable for UAV dynamics.
@@ IV-C3 Interference Management
The system accounts for inter-UAVs interference when computing SINR values. Each UAV’s transmission contributes to the interference perceived by users connected to other UAVs. This coupling is reflected in the environment dynamics and achievable rate calculation. The mechanism encourages UAV agents to learn cooperative spatial and power control strategies that balance coverage maximization and energy minimization.
@@ V-A Reward-based Analysis
Fig. 3 illustrates the convergence of the proposed MADDPG framework. The x-axis represents the Training Episodes (up to 25,000), and the y-axis is the normalized average episodic reward (per step). The average reward rises sharply during the initial exploration phase and stabilizes around 0.032. It exhibits a clear stepped improvement around the 5,000-episode mark, which coincides with the learning-rate decay. This jump is followed by a brief period of high-variance fine-tuning. The policy then settles into its final converged state, and the stable plateau from roughly 8,000 episodes onward confirms robust convergence.
@@
As shown in Fig. 4, we compare the instantaneous reward per step of the learned MADDPG policy against three baselines: KNN-Fixed, QoS-Aware Greedy, and Random. The results demonstrate that the MADDPG policy (red line) consistently maintains the highest average reward across all timesteps.
Specifically, the QoS-Aware Greedy policy (green line) shows competitive performance in the early steps but fluctuates significantly. Although it achieves high instantaneous coverage, its myopic focus on immediate SINR leads to excessive power consumption and inefficient trajectories, thereby reducing the total reward. The KNN-Fixed method (blue line) performs consistently but lacks the adaptability to dynamic traffic changes. In contrast, the MADDPG agents, through cooperative learning, achieve a superior balance between spatial positioning and energy efficiency, enabling them to sustain higher rewards throughout the episode. The Random policy serves as a lower bound, quickly collapsing toward zero reward.
Figure 4: Comparison of instantaneous reward per step among four policies, averaged over 15 testing episodes.
@@ V-B Coverage Performance Evaluation
Figure 5: Comparison of Total Coverage Ratio per Episode.
Fig. 5 illustrates the total coverage ratio per episode. The QoS-Aware Greedy policy achieves the highest instantaneous coverage ($78.42\%$) in most episodes by aggressively directing UAVs toward signal peaks. However, our proposed MADDPG framework achieves a highly comparable coverage rate ($78.37\%$), with a negligible difference of approx. $0.05\%$. This indicates that MADDPG successfully learns to identify and cover critical user clusters almost as effectively as an exhaustive greedy search.
Both algorithms significantly outperform KNN-Fixed ($76.20\%$) and Random, confirming the necessity of intelligent trajectory planning in dynamic sleeping cell scenarios. The performance gap between MADDPG/Greedy and KNN highlights that simply moving to the nearest user (KNN) often leads to suboptimal positioning that neglects cluster-edge users, whereas optimization-based approaches provide better spatial distribution.
V-C Energy Performance Evaluation
Table I presents the critical trade-off between energy and performance. The total energy is decomposed into UAV operational energy ($E_{\text{UAV}}$) and active cell energy ($E_{\text{cell}}$).